# Equivalence Checking of Quantum Circuits via Intermediary Matrix Product Operator

### v1

@@ I Introduction
In our proposed method, an MPO representing $I_{n}$ is positioned between the two circuits, and gates from $G$ and $G^{\prime\dagger}$ are contracted into the MPO using an alternating scheme, $G\rightarrow I_{n}\leftarrow G^{\prime\dagger}$, to exploit the desired equivalence property. This process resembles a double-sided Time Evolving Block Decimation (TEBD) algorithm [8, 9, 10], where gates are applied in a two-qubit temporal zone to neighboring qubit pairs, enhancing computational efficiency and minimizing the growth of bond dimensions—especially when the circuits are equivalent or near-equivalent. Once all gates are contracted into the intermediary MPO, the equivalence of $G$ and $G^{\prime}$ is evaluated by calculating the Frobenius norm of the resulting MPO $GG^{\prime\dagger}$, based on its trace.
@@ II Preliminaries
(b) Contraction of a rank-2 tensor $M_{ij}$ with a rank-1 tensor $v_{j}$ along the shared $j$ index (Matrix-vector multiplication, Eq. (1))
(c) Reshape of a rank-4 tensor into a rank-2 tensor (represented by a dotted bubble around the indices) and vice-versa (represented by a cut through the index, Eq. (2))
(d) SVD of an arbitrary matrix $M$ (represented by a cut through the tensor) with a diamond representing the S matrix to help differentiate it (Eq. (3))
(e) 3-site MPO with labeled indices including dummy indices where $\chi_{i}=1$ as dotted lines (Eq. (4))
Figure 1: Basic Tensor Operations and the MPO structure
In the past decade, tensor networks have emerged as state-of-the-art computational methods for solving problems in quantum many-body physics, from fundamental theoretical challenges to the simulation of quantum circuits. This section aims to provide readers without a tensor network background with a foundation to understand the methods employed in this work. While this introduction is not exhaustive, it serves as a starting point for comprehending the techniques used in our equivalence checking method. For more in-depth treatments of tensor networks, we refer the reader to several resources [8, 11, 12, 9, 10].
Tensors are generalizations of vectors and matrices from linear algebra to multilinear algebra. Computationally, tensors can be viewed as multi-dimensional arrays, with the rank of a tensor referring to the number of dimensions. Scalars, vectors, and matrices are simply rank-0, rank-1, and rank-2 tensors, respectively, while higher-dimensional arrays are referred to as rank-$k$ tensors, where $k\in\mathbb{N}$. Mathematically, a rank-$k$ tensor is indexed by $k$ dimensions and can be visualized as a diagram with $k$ legs extending from a central node, as shown in Fig. 1(a). The symbol used to represent a tensor is arbitrary, but multiple symbols may be employed to highlight different roles in a tensor network.
@@
where $v^{\prime}\in\mathbb{C}^{a}$, and the contraction is performed over the shared index $j$, as visualized in Fig. 1(b).
@@
where $U\in\mathbb{C}^{a\times\chi}$ and $V^{\dagger}\in\mathbb{C}^{\chi\times b}$ are the left- and right-eigenvector matrices, and $S\in\mathbb{R}^{\chi\times\chi}$ is a diagonal matrix of singular values, with $\chi\leq\min(a,b)$. This process, illustrated in Fig. 1(d), is used extensively in tensor network methods. The total number of singular values, $\chi$, is referred to as the bond dimension, which represents the amount of information retained between the two parts of the tensor. Truncating $\chi$ based on a threshold $s_{\text{max}}$ yields an approximate decomposition. This method is frequently used on higher-rank tensors, which are reshaped into matrices, decomposed using the SVD, and then contracted or reshaped back into tensors.
@@
where the physical dimensions $q_{i},q^{\prime}_{i}$ correspond to local transformations at the $i^{\text{th}}$ qubit, and $a_{i-1},a_{i}$ represent the left- and right-bond dimensions. The bond dimensions $\text{dim}(a_{i})=\chi_{i}$ encode the operator’s entanglement capacity, and the MPO structure is shown in Fig. 1(e).
@@ IV MPO-based Equivalence Checking
(c) Application of two-sites of a long-range gate to an MPO
Figure 2: Visualization of MPO operations required for the equivalence checking algorithm
@@ V.2 Long-Range Gates
After constructing $\Theta$, it is reshaped and decomposed using the SVD, similar to the nearest-neighbor case. This process is repeated for all sites until the end of the gate at $\overline{j}$ is reached, at which point the long-range gate MPO has been fully applied to the intermediary MPO. This is visualized in Fig. 2(c).
@@ VI Application Strategy
(b) This shows the equivalent DAG representation of the above circuit. The transition edges represent the qubit it corresponds to such that this representation can help easily identify zones and the order of tensors to apply.
Figure 3: Visualization of the MPO-based equivalence checking algorithm in quantum circuit and DAG form
@@ VI.2 Combined Sweep
In more detail, we perform a two-site spatial sweep across the intermediary MPO $W$, first for even-indexed qubits and then for odd-indexed qubits. At each spatial zone $(i,i+1)$, we apply tensorized gates from both circuits, $G$ and $G^{\prime\dagger}$.
Using a Directed Acyclic Graph (DAG) representation of each quantum circuit [38, 39], we identify the gates within the temporal zone for qubits $(i,i+1)$, allowing us to apply multiple gates at once. We traverse the DAG edges corresponding to these qubits, sequentially applying gates until we encounter a node connected to another qubit outside of the current zone—indicating the end of the temporal zone. As gates are applied, we remove the corresponding nodes from the DAG, and the same method is repeated for later temporal zones in the same spatial zone. Fig. 3 shows an example circuit with its DAG representation, highlighting the spatial and temporal zones.
Once all gates are applied from a given zone, the resulting tensor $\Theta=G_{i,i+1}W_{i}W_{i+1}G^{\prime\dagger}_{i,i+1}$ is decomposed using the SVD, reshaped back into two rank-4 tensors, and used to update the intermediary MPO. The overall even/odd sweep continues across the MPO, and this process is repeated until all gates are applied, i.e., until all nodes have been removed from the DAG.
If a kQG with $k>2$ or a long-range gate is encountered at any spatial zone, we interrupt the even/odd sweep to perform MPO-MPO multiplication. A subsweep is initiated across neighboring sites $(i,i+1)\rightarrow(i+1,i+2)$, contracting both the relevant tensors of the gate MPO and continuing to apply the temporal zone deeper into each circuit.
@@ VI.3 Evaluating $GG^{\prime\dagger}$
$\langle e^{-i\theta}I_{n},W\rangle_{F}=\text{Tr}[W],$
@@
Unlike quantum states, operators do not inherently require normalization, so this trace may not evaluate to 1. However, the identity matrix is the only unitary operator that satisfies $\text{Tr}[I_{n}]=2^{n}$, allowing us to check equivalence according to:
$\frac{\text{Tr}[W]}{2^{n}}\approx 1\Longleftrightarrow G=G^{\prime},$ (16)
where $\frac{\text{Tr}[W]}{2^{n}}\approx 1$ is checked within a tolerance range $1-\epsilon\leq\frac{\text{Tr}[W]}{2^{n}}\leq 1+\epsilon$ for some tolerance $10\leq\epsilon<$ 1.
@@
An alternative way to evaluate $GG^{\prime\dagger}$ is by checking whether the bond dimensions $\chi_{i}=1$ for all $i$. However, in practice, we often observe that bond dimensions can increase slightly, especially for larger systems, even when the singular values are close to zero. While it is possible to perform a final SVD sweep across the entire MPO to reduce these bond dimensions before evaluation, this approach is computationally less efficient than directly calculating the trace.
@@ VII Results
To evaluate the performance and scalability of the proposed MPO-based equivalence checking method, we implemented a Python prototype using numpy [40] for tensor representations and opt-einsum [41] for efficient tensor operations. The performance of this method is compared against the established DD- and ZX-based methods, implemented in C++ as part of MQT-QCEC [26], which is included in the Munich Quantum Toolkit (MQT) [42].
For benchmarking, we generated parameterized two-local quantum circuits using IBM’s qiskit [43]. Each circuit consists of a layer of single-qubit rotation gates, $R_{x}(\theta)$, followed by an entanglement layer of two-qubit rotation gates, $R_{zz}(\theta)$, and another layer of $R_{x}(\theta)$ gates. The structure of each circuit is as follows:
[…]
These circuits are generated as square $n\times n$ circuits, meaning that for $n$ qubits, there are $n$ repetitions of this rotation structure. Each rotation angle, $\theta$, is randomly (according to a uniform distribution) selected from the range $[-\pi,\pi]$, with each block generated independently.
To evaluate the effectiveness of the MPO-based method, we considered three classes of entanglement patterns, each with varying numbers of long-range gates:
@@
    Each qubit interacts only with its nearest neighbors.
@@
    Shifted-Circular Alternating (SCA)
    Entanglement
    Each layer contains a single long-range gate connecting the first and last qubits, shifting between layers [44].
@@
    Each qubit interacts with every other qubit, resulting in many long-range gates.
@@
VII.1 Error Injection and SVD Behavior
The SVD threshold is critical in the performance of tensor network methods, and understanding its behavior is essential before analyzing the scalability of the proposed MPO-based method. To investigate this, we tested three SVD thresholds, $s_{\text{max}}$: a low threshold ($10^{-6}$), a medium threshold ($10^{-3}$), and a high threshold ($10^{-1}$). Since the equivalence condition $G=G^{\prime}$ holds only when the intermediary MPO exactly equals the identity and non-equivalent cases have infinite variations, we expect flexibility in choosing $s_{\text{max}}$ without risking false equivalencies.
@@ VII.1 Error Injection and SVD Behavior
VII.1.1 Missing Gate Errors
@@
VII.1.2 Rotation Angle Errors
@@ VII.1.3 Permutation Errors
VII.2 Scaling of Equivalent Circuits
This section evaluates the scaling behavior of different methods for verifying the equivalence of circuits, repeated for increasing values of $n$, to assess performance with circuit size. We compare the proposed MPO-based method against the DD- and ZX-based methods, with the results shown in Fig. 6.

### v2

@@ I Introduction
In our proposed method, an MPO representing $I_{n}$ is positioned between the two circuits, and gates from $G$ and $G^{\prime\dagger}$ are contracted into the MPO using an alternating scheme, $G\rightarrow I_{n}\leftarrow G^{\prime\dagger}$, to exploit the desired equivalence property. This process resembles a double-sided time-evolving block decimation (TEBD) algorithm [8, 9, 10], where gates are applied in a two-qubit temporal zone to neighboring qubit pairs, enhancing computational efficiency and minimizing the growth of bond dimensions—especially when the circuits are equivalent or near-equivalent. Once all gates are contracted into the intermediary MPO, the equivalence of $G$ and $G^{\prime}$ is evaluated by calculating the Frobenius norm of the resulting MPO $GG^{\prime\dagger}$, based on its trace.
@@ II Preliminaries
(b) Contraction of a rank-2 tensor $M_{ij}$ with a rank-1 tensor $v_{j}$ (Matrix-vector multiplication, Eq. (1))
(c) Reshape of a rank-4 tensor into a rank-2 tensor and vice versa (Eq. (2))
(d) SVD of an arbitrary matrix $M$ (Eq. (3)); the diamond denotes the $S$ matrix
(e) A 3-site MPO with labeled indices, including $\chi_{i}=1$ dotted lines (Eq. (4))
Figure 1: Basic tensor operations and the MPO structure. Each panel (a)–(e) illustrates one fundamental concept or structure.
In the past decade, tensor networks have emerged as state-of-the-art computational methods for solving problems in quantum many-body physics, from fundamental theoretical challenges to the simulation of quantum circuits. This section aims to provide readers without a tensor network background with a foundation to understand the methods employed in this work. While this introduction is not exhaustive, it serves as a starting point for comprehending the techniques used in our equivalence checking method. For more in-depth treatments of tensor networks, we refer the reader to several resources [8, 11, 12, 9, 10]. Additionally, a visualization of the operations used throughout this work are found in Fig. 1.
Tensors are generalizations of vectors and matrices from linear algebra to multilinear algebra. Computationally, tensors can be viewed as multi-dimensional arrays, with the rank of a tensor referring to the number of dimensions. Scalars, vectors, and matrices are simply rank-0, rank-1, and rank-2 tensors, respectively, while higher-dimensional arrays are referred to as rank-$k$ tensors, where $k\in\mathbb{N}$. Mathematically, a rank-$k$ tensor is indexed by $k$ dimensions and can be visualized as a diagram with $k$ legs extending from a central node. The symbol used to represent a tensor is arbitrary, but multiple symbols may be employed to highlight different roles in a tensor network.
@@
where $v^{\prime}\in\mathbb{C}^{a}$, and the contraction is performed over the shared index $j$.
@@
where $U\in\mathbb{C}^{a\times\chi}$ and $V^{\dagger}\in\mathbb{C}^{\chi\times b}$ are the left- and right-eigenvector matrices, and $S\in\mathbb{R}^{\chi\times\chi}$ is a diagonal matrix of singular values, with $\chi\leq\min(a,b)$. This process is used extensively in tensor network methods. The total number of singular values, $\chi$, is referred to as the bond dimension, which represents the amount of information retained between the two parts of the tensor. Truncating $\chi$ based on a threshold $s_{\text{max}}$ yields an approximate decomposition. This method is frequently used on higher-rank tensors, which are reshaped into matrices, decomposed using the SVD, and then contracted or reshaped back into tensors.
@@
where the physical dimensions $q_{i},q^{\prime}_{i}$ correspond to local transformations at the $i^{\text{th}}$ qubit, and $a_{i-1},a_{i}$ represent the left- and right-bond dimensions. The bond dimensions $\text{dim}(a_{i})=\chi_{i}$ encode the operator’s entanglement capacity.
@@ IV MPO-based Equivalence Checking
(c) Application of two sites of a long-range gate to an MPO
Figure 2: Visualization of MPO operations required for the equivalence checking algorithm.
@@ V.2 Long-range gates
After constructing $\Theta$, it is reshaped and decomposed using the SVD, similar to the nearest-neighbor case. This process is repeated for all sites until the end of the gate at $\overline{j}$ is reached, at which point the long-range gate MPO has been fully applied to the intermediary MPO. This is visualized in Fig. Fig. 2(c).
@@ VI Application Strategy
(b) This shows the equivalent DAG representation of the above circuit. The transition edges represent the qubit it corresponds to, allowing easy identification of zones and the order of tensors to apply.
Figure 3: Visualization of the MPO-based equivalence checking algorithm in both quantum circuit (a) and DAG (b) form.
@@ VI.2 Combined sweep
In this method, we perform a two-site spatial sweep across the intermediary MPO $W$, alternating between even-indexed and odd-indexed qubits. At each spatial zone $(i,i+1)$, we apply tensorized gates from both circuits, $G$ and $G^{\prime\dagger}$, in the order defined by the directed acyclic graph (DAG) representation of each circuit.
The DAG representation provides a structured way to identify temporal zones, where we traverse the edges corresponding to qubits $(i,i+1)$ and sequentially apply gates until encountering a node connected to another qubit outside the current spatial zone. This traversal marks the end of the temporal zone for $(i,i+1)$. Nodes corresponding to applied gates are removed from the DAG, and the process is repeated for subsequent spatial zones. As shown in Fig. 3, the DAG representation helps determine the temporal zones and their associated gate sequences.
Differences in the DAG structures of $G$ and $G^{\prime}$ can introduce mismatches in the temporal zones, particularly when long-range gates or differences in qubit connectivity are present. For example, while the transpilation process for $G^{\prime}$ may assume all-to-all qubit connectivity, $G$ may retain a specific structure such as a linear nearest-neighbor structure. This mismatch results in temporal zones in $G^{\prime}$ that may not perfectly align with those in $G$. In such cases, long-range gates in either circuit require recursive treatment.
More precisely, to ensure efficient operator cancellations even in cases of long-range gates, the method introduces an MPO-MPO contraction step when a long-range gate or $k$-qubit gate ($k>2$) is encountered. Here, the overall even/odd sweep is interrupted to handle the long-range interaction by using a subsweep along the sites of the gate MPO. As each set of neighboring sites in the gate MPO are contracted with the intermediary MPO, the temporal zone is applied from deeper into each circuit. This approach minimizes the computational cost associated with handling mismatched DAG structures while preserving the operator cancellations enabled by local temporal zones. By addressing potential mismatches in the DAG structures of $G$ and $G^{\prime}$ through recursive subzone processing, this method ensures robustness and efficiency even in the presence of long-range gates or varying circuit connectivity patterns.
Once all gates are applied for a given spatial zone, the resulting tensor $\Theta=G_{i,i+1}W_{i}W_{i+1}G^{\prime\dagger}_{i,i+1}$ is decomposed using the SVD, reshaped back into two rank-4 tensors, and used to update the intermediary MPO. The even/odd sweep continues across the MPO until all gates have been applied and all nodes removed from the DAG.
@@ VI.3 Evaluating $GG^{\prime\dagger}$
$\langle e^{-i\theta}I_{n},W\rangle_{F}=e^{i\theta}\ \text{Tr}[W],$
@@
Unlike quantum states, operators do not inherently require normalization; thus, the trace of $W$ may not be 1. However, for an $n$-qubit identity operator $I_{n}$, we have $\mathrm{Tr}[I_{n}]=2^{n}$, making $I_{n}$ (up to a global phase) the unique unitary achieving that maximum trace. Consequently, we can check equivalence via
$\biggl{|}\,e^{i\theta}\,\frac{\mathrm{Tr}[W]}{2^{n}}\biggr{|}\approx 1\quad\Longleftrightarrow\quad G=G^{\prime},$
which in practice leads to the condition
[…]
where $\epsilon$ ($10\leq\epsilon<$ 1) defines a numerical tolerance such that
[…]
Furthermore, the global phase does not affect $\tfrac{\mathrm{Tr}[W]}{2^{n}}$ because
$\biggl{|}\,e^{i\theta}\,\frac{\mathrm{Tr}[W]}{2^{n}}\biggr{|}\;=\;|\,e^{i\theta}\,|\,\biggl{|}\frac{\mathrm{Tr}[W]}{2^{n}}\biggr{|}\;=\;\biggl{|}\frac{\mathrm{Tr}[W]}{2^{n}}\biggr{|}.$
@@
An alternative way to verify $GG^{\prime\dagger}$ is to first check whether all bond dimensions $\chi_{i}=1$ for all $i$, then check if each local matrix is the identity. However, in practice, small numerical imprecisions often cause slight increases in bond dimensions, especially for larger systems, even when the result is the identity. A final SVD sweep across the entire MPO can reduce these extra dimensions, but it is computationally less efficient than directly calculating $\mathrm{Tr}[GG^{\prime\dagger}]$. Moreover, $\chi_{i}=1$ could also mean that the operator factorizes into arbitrary local operators, making this criterion ambiguous. By contrast, computing the trace immediately provides a clear global check of equivalence that is both more robust and efficient in most practical settings.
@@ VII Results
To evaluate the performance and scalability of the proposed MPO-based equivalence checking method, we implemented a Python prototype using numpy [38] for tensor representations and opt-einsum [39] for efficient tensor operations. This implementation is part of the Munich Quantum Toolkit (MQT) [40] and is available in the YAQS package found at [41], as well as integrated as a tool in QCEC [26]. The performance of this method is compared against the established DD- and ZX-based methods found in the QCEC package. All tests were executed on an i7-1165G7 (2.80 GHz).
For benchmarking, we generated parameterized two-local quantum circuits using IBM’s qiskit [42]. Each circuit consists of a block made of a layer of single-qubit rotation gates, $R_{x}$, followed by an entanglement layer of two-qubit rotation gates $R_{zz}$ (with various entanglement patterns). This is then repeated before a final layer of $R_{x}$ gates.
The tested circuits are then generated as square $n\times n$ circuits, meaning that for $n$ qubits, there are $n$ repetitions of this rotation structure. Each rotation angle, $\theta_{i}$, is randomly (according to a uniform distribution) selected from the range $[-\pi,\pi]$, with each block generated independently such that the structures do not repeat. This is visualized in Fig. 5(a).
To evaluate the performance of the MPO-based method, we consider three classes of entanglement patterns, each distinguished by how many long-range gates they contain:
@@
    Each qubit interacts only with its nearest neighbors. In each layer, the $R_{zz}$ gates follow a staircase pattern, connecting $(q_{i},q_{i+1})$ for $i=1,\dots,n$.. Hence, there are no long-range gates. This is shown in Fig. 5(b)
@@
    Shifted-circular alternating (SCA) entanglement
    This pattern extends the linear circuit by adding a single long-range gate $(q_{1},q_{n})$ within each layer. The order of gates in a layer shifts so that the long-range gate alternates its position each time [43]. For example, in one layer the entangling gates contains a long-range gate $(q_{1},q_{n})$ followed by $(q_{i},q_{i+1})$ for $i=1,\dots,n$. In the next layer, this long-range gate shifts such that the circuit has a gate $(q_{1},q_{2})$, then the long-range $(q_{1},q_{n})$, followed by $(q_{i},q_{i+1})$ for $i=2,\dots,n$. The first repetition block is shown in Fig. 5(c) as an example.
@@
    Each qubit interacts with every other qubit in each layer, resulting in many long-range gates. Specifically, the $R_{zz}$ gates connect all pairs $(q_{i},q_{j})$ for $i,j=1,\dots,n$ where $i\neq j$ creating both a larger total number of gates and extensive long-range entanglement. This is shown in Fig. 5(d).
@@
[…]
(a) This visualizes the general circuit structure used in all experimental setups. This structure is repeated $n$ times, each time with different random rotation angles, using the entanglement layers below.
[…]
(b) Linear entanglement layer
[…]
(c) Shifted-circular alternating (SCA) entanglement layer
[…]
(d) Full entanglement layer
Figure 5: This figure shows a single repetition of each circuit structures used in the experimental setup. For $n$ qubits, each structure is repeated $n$ times. All angles are randomly selected for each gate according to a uniform distribution $[-\pi,\pi]$. Each repetition is generated independently with independently-generated angles.
VII.1 SVD behavior
The SVD threshold $s_{\text{max}}$ plays a critical role in the performance of tensor network methods, as it directly controls the truncation of singular values, and thus, the growth of the bond dimensions of the intermediary MPO. Understanding its effect is essential before analyzing the scalability of the proposed MPO-based equivalence-checking method.
To investigate this, we generate 10 $5\times 5$ circuits $G$ of each entanglement type (linear, SCA, and full), and transpile them into new circuits $G^{\prime}$. We then check their equivalence using 60 logarithmically spaced SVD thresholds $s_{\text{max}}$ in the range $[10^{-6},1]$.
For each $s_{\text{max}}$, we compute the deviation from the ideal equivalence condition as:
$\epsilon(s_{\text{max}})=1-\frac{\text{Tr}[W]}{2^{n}}.$
This $\epsilon(s_{\text{max}})$ measures the inherent numerical error generated from the truncation in the method itself. This can be understood as the numerical tolerance required in the trace to assert equivalence between the two circuits $G$ and $G^{\prime}$ for a given SVD threshold such that $\epsilon<\epsilon(s_{\text{max}})$ must be true to guarantee correctness. The average over 10 circuits of each entanglement type is plotted in Fig. 6.
The plot illustrates how $\epsilon(s_{\text{max}})$ behaves for the three entanglement patterns, highlighting the threshold beyond which equivalence can no longer be reliably asserted. For instance, circuits with full entanglement exhibit significant sensitivity to the SVD threshold, continuously introducing more error as the threshold increases. In contrast, the SCA circuit with fewer long-range gates has a hard-cut from roughly machine error to a continuous curve around $s_{\text{max}}\approx 10^{-2}$. Notably, linear circuits are robust for any value $s_{\text{max}}<1$. Due to the cancellations that occur in these benchmarks, the intermediary MPO will almost exclusively be close to the identity. As soon as the intermediary MPO is not consistently close, e.g., in the case of many long-range gates, then this value must be chosen with care.
Figure 6: This figure shows the deviation $\epsilon(s_{\text{max}})=1-\frac{\text{Tr}[W]}{2^{n}}$ caused by various SVD thresholds for each circuit type. This was tested for 60 logarithmically-spaced SVD thresholds $s_{\text{max}}$ for 10 circuits of each type, where the average is plotted. The shaded area shows one standard deviation from the average.
@@ VII.1 SVD behavior
VII.2 Error injection
To investigate this, we tested three SVD thresholds, $s_{\text{max}}$: a low threshold ($10^{-6}$), a medium threshold ($10^{-3}$), and a high threshold ($10^{-1}$). Since the equivalence condition $G=G^{\prime}$ holds only when the intermediary MPO exactly equals the identity and non-equivalent cases have infinite variations, we expect flexibility in choosing $s_{\text{max}}$ without risking false equivalencies.
@@ VII.2 Error injection
VII.2.1 Missing gate errors
@@
VII.2.2 Rotation angle errors
@@ VII.2.3 Permutation errors
VII.3 Scaling of equivalent circuits
This section evaluates the scaling behavior of different methods for verifying the equivalence of circuits, repeated for increasing values of $n$, to assess performance with circuit size. We compare the proposed MPO-based method against the DD- and ZX-based methods, with the results shown in Fig. 8. All tests have an SVD threshold set such that we guarantee equivalence with a tolerance $\epsilon=10^{-13}$, i.e., linear $s_{\text{max}}=10^{-1}$, SCA $s_{\text{max}}=10^{-3}$, full $s_{\text{max}}=10^{-6}$.